OpenStudy (anonymous):

Is the expression x3•x3•x3 equivalent to x3•3•3? Why or why not? Explain your reasoning

OpenStudy (studio-games):

Ok lemme see.

OpenStudy (welshfella):

I meant x^3 * x^3 * x^3 = x^(3+3+3) = x^9

OpenStudy (welshfella):

and in the second expression you only have one x^3

OpenStudy (welshfella):

- and you are multiply it by 3 * 3